# Impact of chaotic magnetic field on mass-radius relation of rotating neutron stars

### v3

@@ Abstract
Observations reveal that magnetic fields on neutron stars (NSs) are in the range of $10^{8-15}$ G. Apart from being celestial bodies, NSs are normally rotating. In this work, we study the impact of a chaotic magnetic field on the mass-radius relation of the rotating NSs. We employ an equation of state of NSs with the nuclei in the crust and hyperons in the core. We use Hartle-Thorne formalism as an approximation of the rotating NSs. For the magnetic field ansatz, we use the one coupled to the energy density. We find that the magnetic field can decrease radius of NS. In contrast, the increment of the magnetic field increases the compactness and deformation of rotating NSs.
@@ 3 Results and Discussion
Fig. 1 presents the mass-radius relationship of NSs in a static configuration. The presence of a chaotic magnetic field reduces the radius of NSs, as indicated by the leftward shift of the curves with increasing magnetic field strength, suggesting that a chaotic magnetic field contributes to greater compactness of the stars. Compactness, defined as the mass-to-radius ratio, is directly influenced by this field. Notably, the radius reduction is pronounced in the case where $B_{0}=3\times 10^{18}$ G and $B_{surf}=10^{15}$ G, indicating that the radius decrease is non-linear with respect to the increase in the chaotic magnetic field. However, our results diverge from those reported by the authors in Ref. [12], where the chaotic magnetic field was found to increase the radius of NSs. This discrepancy may be attributed to the different EOS employed in their study. Furthermore, it is worth noting that as the magnetic field strength increases, the mass-radius curve begins to resemble that of quark stars. For more detailed discussions on quark stars, please refer to Refs. [23, 24, 25].
@@ 4 Conclusion
In this study, we calculated the mass and radius of NSs under the influence of a chaotic magnetic field, enabling us to obtain the mass-radius relationship for these stars. Our findings indicate that the magnetic field decreases the radius of NSs. Additionally, while the chaotic magnetic field contributes to a reduction in radius, it simultaneously enhances the compactness and deformation of rotating NSs, with significant deformation occurs at lower mass values.

### v4

@@ Abstract
Observations reveal that magnetic fields on neutron stars (NSs) are in the range of $10^{8-15}$ G. Apart from being celestial bodies, NSs are normally rotating. In this work, we study the impact of a chaotic magnetic field on the mass-radius relation of the rotating NSs. We employ an equation of state of NSs with the nuclei in the crust and hyperons in the core. We use Hartle-Thorne formalism as an approximation of the rotating NSs. For the magnetic field ansatz, we use the one coupled to the energy density. We find that the magnetic field can decrease radius of NS. NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. In contrast, the increment of the magnetic field can increase the compactness and deformation of rotating NSs.
@@ 3 Results and Discussion
Fig. 1 presents the mass-radius relationship of NSs in a static configuration. The presence of a chaotic magnetic field can reduce the radius of NSs, as indicated by the leftward shift of the curves with increasing magnetic field strength, suggesting that a chaotic magnetic field can contribute to greater compactness of the stars. Compactness, defined as the mass-to-radius ratio, is directly influenced by this field. Notably, the radius reduction is pronounced in the case where $B_{0}=3\times 10^{18}$ G and $B_{surf}=10^{15}$ G, indicating that the radius decrease is non-linear with respect to the increase in the chaotic magnetic field. Furthermore, NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. Our findings, however, differ from those presented by Lopes et al. [12], who reported that chaotic magnetic fields increase the radius of NSs and allow for the formation of NSs with a higher maximum mass when the chaotic magnetic field is stronger. This discrepancy may be attributed to the different EOS employed in their study. Additionally, it is worth noting that as the magnetic field strength increases, the mass-radius curve begins to resemble that of quark stars. For more detailed discussions on quark stars, please refer to Refs. [23, 24, 25].
@@ 4 Conclusion
In this study, we calculated the mass and radius of NSs under the influence of a chaotic magnetic field, enabling us to obtain the mass-radius relationship for these stars. Our findings indicate that the magnetic field can decrease the radius of NSs. NSs formed with stronger chaotic magnetic fields exhibit a lower maximum mass compared to those formed with weaker chaotic magnetic fields. Additionally, while the chaotic magnetic field can contribute to a reduction in radius, it can simultaneously enhance the compactness and deformation of rotating NSs, with significant deformation occurs at lower mass values.